# SOUS VIDE: Cooking Visual Drone Navigation Policies in a Gaussian Splatting Vacuum

### v1

@@ V SV-Net Policy Architecture
Our policy architecture, SV-Net, runs on an Orin Nano onboard the drone at 20 Hz and is illustrated in Fig. 4. The policy relies exclusively on data directly observable by the drone: (i) images from the onboard camera and (ii) state estimates of vertical position, translational velocities, and quaternion orientation ($p^{k}_{z\mathcal{W}},\bm{v}^{k}_{\mathcal{W}},\bm{q}^{k}_{\mathcal{BW}}$), derived from an EKF. These estimates are typically obtained on a drone by fusing data from an IMU, a downward-facing time-of-flight sensor, and an optical flow sensor—all of which are mature, small-footprint, and plug-and-play units commonly available on hobby-grade drones. The policy outputs thrust and body rate commands $\bm{u}^{k}=(f_{th},\bm{\omega})$, a standard control mode used by FPV drone racing pilots.
SV-Net comprises three components: a feature extractor, a history network and a command network. The architecture uses SqueezeNet [12] as a vision encoder, augmenting its output with estimated height and orientation before passing it through an MLP to create a pose-aware feature extractor. The history network uses a sliding T-step window of the observable state to generate a latent vector encoding the evolving flight dynamics of the drone at that instant, a technique called Rapid Motor Adaptation (RMA). The policy ingests the latent vector to adapt its output to current flight conditions. The command network combines the outputs of the feature extractor and history network with the observable states and a user-defined objective $\mathcal{O}^{k}$. This objective encodes the first and last waypoints of the trajectory and the total trajectory time. This remains constant at deployment except for when we use a single policy to encode multiple trajectories.
@@
Figure 4: The data pre-processing step used in SV-Net policy. The blue hue of the processed image, $\mathcal{I}^{k}_{proc}$, illustrates the color normalization in contrast to the raw image, $\mathcal{I}^{k}_{proc}$, with both displayed using a [0,1] visualization range.
The policy inputs are derived from onboard sensor data using the pre-processing outlined in Fig. 4, where raw images are scaled and normalized, and observable EKF data is processed through the history transform defined by:
$\begin{split}\delta t^{k-1}&=t^{k}-t^{k-1}\\ \delta\bm{p}_{\mathcal{W}}^{k-1}&=\delta t^{k-1}\cdot\bm{v}^{k}\\ \delta\bm{v}_{\mathcal{W}}^{k-1}&=\bm{v}^{k}-\bm{v}^{k-1}\\ \delta\bm{q}_{\mathcal{BW}}^{k-1}&=\bm{q}_{\mathcal{WB}}^{k}\cdot\bm{q}_{\mathcal{BW}}^{k-1}.\end{split}$ (3)
Note that we use velocity to back out a change in position for the RMA module. We do not use positional information from VIO or external motion capture for this.
Figure 5: Clockwise from top left: 1) Flight test objectives shown as a 3D projection onto the Gaussian Splat and in real-world FPV. 2) Drone hardware setup with frames $(\mathcal{W},\mathcal{B},\mathcal{C})$. We use a PixRacer Pro for low-level control and an Orin Nano for policy execution. Sensing is done using the PixRacer’s IMU, an ARK Flow sensor, motion capture markers and the D435’s monocular camera. 3) 3D plots depicting real-world performance of the policies in Section VI-A.
To train the policy, we first pre-train the RMA network on the sampled drone parameters. Once trained, the RMA network is frozen, and the remaining components of the policy are trained end-to-end (including the SqueezeNet image encoder) to predict body rate commands. In hardware testing, we find the best performance is achieved using the output of the second-to-last layer, not the explicit estimate of $\bm{\theta}$, as input into the command network MLP.
@@ V-A Analysis of RMA Module
A property of (1) is that if we allow the drone parameters, $k_{th}$ and $m$, to be variables that can be adjusted online, we can use them to compensate for a wide range of model inaccuracies that are not limited to the thrust and weight of the drone. For simplicity, let $c=\frac{k_{th}}{m}$. Given an additional force vector $\bm{f}_{add}$ in the world frame to account for model inaccuracies from the thrust coefficient, mass and also for any external forces like aerodynamic drag or ground effect, we can compute an equivalent $\hat{c}$ in an augmented form of the velocity equation in (1). We start by doing
$\begin{split}-g\bm{z}_{\mathcal{W}}+\hat{c}f_{th}\bm{z}_{\mathcal{B}}&=-g\bm{z}_{\mathcal{W}}+cf_{th}\bm{z}_{\mathcal{B}}+\bm{f}_{add}\end{split}$ (4)
and then taking the least-squares estimate of $\hat{c}$, we get:
$\begin{split}\hat{c}&\approx\frac{c+\bm{z}_{\mathcal{B}}^{T}\bm{f}_{add}}{\hat{f}_{th}}\end{split}.$ (5)
@@ VI Experiments
In this section, we evaluate our SOUS VIDE policies across three fronts: efficacy of the proposed policy architecture, robustness to dynamic and visual disturbances, and generalization to novel scenarios. We demonstrate that the SV-Net policy, equipped with the RMA module, achieves state-of-the-art performance in zero-shot sim-to-real transfer. We emphasize that in all our experiments, the policy does not get to observe lateral position $(p_{x\mathcal{W}},p_{y\mathcal{W}})$, e.g., from a motion capture system, or an onboard VIO system. We do obtain vertical position $p_{z\mathcal{W}}$ from the onboard time-of-flight sensor.
We perform all experiments using a quadrotor drone equipped with a PixRacer Pro for low-level body-rate tracking control and an Orin Nano for policy execution, as shown in Fig. 5. The onboard sensing suite includes an IMU, ARK Flow sensor, and a monocular camera. The motion capture markers visible in our images and videos are for diagnostics (to plot trajectories and compare against ground truth). To assess performance, we evaluate the following metrics:
  • Success: Completes the intended trajectory. We subdivide this into: 1) no collisions and 2) collisions that recover.
  • Partial Success: Tracks the position, but not orientation.
  • Failure: Fails to finish the trajectory. We subdivide this into: 1) collisions that do not recover 2) drifts off-course.
  • Collision Rate: The number of collisions the drone makes, normalized by the total distance it traveled.
  • Trajectory Tracking Error: The $\ell_{2}$-norm of the drone’s position to the nearest point on the desired trajectory.
  • Proximity Percentile: The percentage of the actual trajectory that is within a distance $r\!=\!0.3m$ of the desired.
These metrics ensure a comprehensive evaluation of trajectory adherence, robustness, and recovery behavior.
Ablation Experiments Success Partial Success Failure
Collision
Rate
Trajectory
Tracking Error
Proximity
Percentile
No Collisions Recovery No Recovery Off-Course (hits/m) (m) (%)
SV no RMA 3 0 1 0 1 0.00 0.646 37.7
SV no pre-train 0 0 1 4 0 0.10 0.392 51.4
SV no latent 1 2 1 1 0 0.05 0.434 42.0
SV-Net (ours) 5 0 0 0 0 0.00 0.173 96.0
TABLE I: Table comparing the performance of the SV-Net and its architecture ablations
Robustness Experiments Success Partial Success Failure
Collision
Rate
Trajectory
Tracking Error
Proximity
Percentile
No Collisions Recovery No Recovery Off-Course (hits/m) (m) (%)
Baseline 5 0 0 0 0 0.00 0.173 96.0
Lighting 4 0 1 0 0 0.00 0.237 64.2
Dynamic 5 0 0 0 0 0.00 0.188 86.3
Static 0 0 4 0 1 0.06 0.488 25.3
Payload 5 0 0 0 0 0.00 0.202 80.1
Wind 5 0 0 0 0 0.00 0.197 81.7
TABLE II: Table presenting SV-Net performance under visual and dynamic disturbances.
@@ VI-A Policy Architecture Ablations
We evaluated our main policy versus three ablations on a 15-second trajectory that guides the drone through a gate and under a ladder before ending facing a monitor. The desired trajectory is visualized in the GSplat in Fig. 5, along with flights from each policy ablation. All policies were trained on the same expert MPC dataset using PyTorch, the Adam optimizer (learning rate 1e-4), for approximately 12 hours on a desktop machine (i9-13900K, RTX 4090, 64GB RAM). The policy ablations are:
@@
  • SV no RMA: A minimal variant comprising only the feature extractor and command network. This serves as our closest approximation to a zero-shot equivalent of the few-shot policy described in [14].
@@
As shown in Table II, we found that SV-Net outperformed all other architectures, achieving a success rate of $100\%$ and a trajectory tracking error of 0.173m, more than doubling the performance of SV no RMA. Pre-training the history network proved essential because (we hypothesize) it allows the network to adapt to significant changes in the flight dynamics due to rotor effects close to obstacles and due to significant effects from battery drain throughout a trajectory and across multiple trajectories. Also, we suspect that using the 2nd-to-last layer instead of the last layer of the history network enhances performance as the higher dimensionality of the latent code outweighs the minimal information loss from excluding the final layer. Lastly we observe, albeit anecdotally, that when training on significantly larger datasets, SV no pre-train’s performance in trajectory tracking converges with SV-Net when tested within FiGS. However, at this scale, the associated training times become impractical.
@@ VI-B Robustness Experiments
We execute the same SV-Net policy as in Section VI-A across five experiments, each introducing distinct disturbances, as shown in Fig. 6:
@@
  • Dynamic: Four people moved dynamically within the field of view along the trajectory.
  • Static: The gate, ladder, and monitor (present at train time) were removed at runtime while the pillars adjacent to the gate were occluded.
  • Payload: A rigid 350g payload ($30\%$ increase in drone weight) was attached.
@@
Figure 6: Visualization of disturbances and the corresponding position and velocity performance of SV-Net. Lighting: illumination reduced by 60%, Dynamic: people moving erratically in the scene, Static: objects from training removed or occluded at runtime, Payload: 30% increase in mass, Wind: 40 m/s gust from leaf blower. SV-Net maintains adequate performance in all cases.
Scenario Success Partial Success Failure
Collision
Rate
Trajectory
Tracking Error
Proximity
Percentile
No Collisions Recovery No Recovery Off-Course (hits/m) (m) (%)
Multi-Objective 15 0 0 0 0 0.00 0.203 83.4
Extended Trajectory 23 4 0 3 0 0.02 0.241 72.5
Cluttered Environment 13 1 0 0 1 0.01 n/a n/a
TABLE III: Table showcasing the performance of SV-Net across three different skills.
The results in Table II show SV-Net consistently demonstrated resilience to dynamic disturbances, payload variations, and wind gusts, maintaining near-baseline performance with only marginal increases in trajectory tracking error. However, lighting changes led to partial failures under extremely low lumens, where distinguishing objects, such as a black monitor blending into the dim background, became challenging. We also note that when lighting dropped below $40\%$ of the original lumens, the policy consistently failed by drifting off-course from the start location. Static scene changes proved to be the most challenging: (i) the policy underflew waypoints and experienced minor collisions with the occluded pillars, (ii) the policy consistently flew through the space where the rungs of the ladder would have been. Despite these difficulties, the policy reliably tracked the overall trajectory shape, recovered from collisions, and successfully reached the final position in 4 out of 5 flights. These results suggest that the policy retains critical scene information that would otherwise have been removed in approaches relying on visual abstractions.
@@ VI-C Skill-Testing Experiments
In our last set of experiments, we trained three different SV-Net policies, one for each of the following scenarios:
  • Multi-Objective: One policy was trained to execute three distinct trajectories within the same scene distinguished by unique objective inputs $O^{k}$ (start and end points, and total trajectory time). Two trajectories used identical positional splines but traversing in opposite directions, while the third followed a climbing orbit.
@@
  • Cluttered Environment: The policy was deployed close to the ground in a highly cluttered workshop with obstacles spaced as close as 1.0 m apart. The workshop has no motion capture system, so no ground truth trajectory data is available. Instead, we visualize the flight using a time-lapse rendering.
Results are reported in Table III with illustration shown in Fig. 7. The multi-objective policy had mixed success, indicating the need for more robust objective encodings in future work. Though it achieved a $100\%$ collision free success rate, we observed significant degradation in one of the three tasks, where the drone consistently under-flew the desired trajectory and over-shot its end point. In the extended trajectory, the policy performed comparably to shorter trajectories, maintaining low collision rates and a Proximity Percentile of $72.5\%$ over 30 attempts. Finally, in the cluttered environment, the policy achieved a $93.3\%$ success rate on a 25s trajectory in between tables in a visually complex environment, highlighting its robustness in practical applications without relying on external pose estimators. Since the environment lacked motion capture, we cannot measure or report Trajectory Tracking Error or Proximity Percentile.
Figure 7: 3D plots depicting performance of the skill-testing policies in Section VI-C. We show a time-lapse trajectory for the cluttered environment as this room does not have a motion capture system.

### v2

@@ V SV-Net Policy Architecture
Our policy architecture, SV-Net, runs on an Orin Nano onboard the drone at 20 Hz. To output thrust and body rate commands, $\bm{u}=(f_{th},\bm{\omega})$, the policy relies solely on onboard data: (i) images from the onboard camera and (ii) height, velocity, and orientation estimates ($p_{z},\bm{v}_{\mathcal{W}},\bm{q}_{\mathcal{BW}}$) provided by an Extended Kalman Filter (EKF), which fuses data from an IMU, a downward-facing time-of-flight sensor, and an optical flow sensor. These inexpensive, compact sensors are common on hobby-grade drones, providing state estimates that, while not pinpoint precise, are useful for control—especially since most height-sensitive applications occur over reasonably level surfaces. Notably, SV-Net performs better with $(p_{z},\bm{v}_{\mathcal{W}})$, even when overflying obstacles, than without them. Beyond serving as direct inputs to SV-Net, these estimates, along with timestamps, are used to compute the history data:
$\begin{split}\delta t^{k-1}=t^{k}-t^{k-1},\qquad\delta\bm{p}_{\mathcal{W}}^{k-1}=\delta t^{k-1}\cdot\bm{v}_{\mathcal{W}}^{k},\\ \delta\bm{v}_{\mathcal{W}}^{k-1}=\bm{v}_{\mathcal{W}}^{k}-\bm{v}_{\mathcal{W}}^{k-1},\qquad\delta\bm{q}_{\mathcal{BW}}^{k-1}=\bm{q}_{\mathcal{WB}}^{k}\cdot\bm{q}_{\mathcal{BW}}^{k-1}.\end{split}$ (3)
We use $\bm{v}_{\mathcal{W}}$ to infer $\delta\bm{p}$ as the drone cannot observe its lateral position. For brevity, we use superscript time indices.
@@
SV-Net comprises three components: a feature extractor, a history network and a command network (Fig. 3). The architecture uses SqueezeNet [12] as a vision encoder, augmenting its output with estimated height and orientation before passing it through an MLP to create a pose-aware feature extractor. The history network, inspired by RMA, uses the sliding time-step window of history data to generate a latent vector encoding the evolving flight dynamics of the drone at that instant. The policy ingests the latent vector to adapt its output to current flight conditions. The command network combines the outputs of the feature extractor and history network with the observable states and an objective vector, $\mathcal{O}^{k}$, which encodes the change in position, initial and final velocity, initial and final orientation (quaternion), and total trajectory time. We use this to facilitate training and deployment across different trajectories when a single SV-Net is encoded with multiple trajectories (Section VI-C).
Figure 4: Clockwise from top left: 1) Desired trajectory in the scene’s GSplat with corresponding real-world First-Person-View (FPV) of key objects. 2) Drone hardware and frames $(\mathcal{W},\mathcal{B},\mathcal{C})$. We use an Orin Nano and PixRacer Pro for control, while sensing is handled by the PixRacer’s IMU, an ARK Flow sensor, and the D435’s monocular camera. Motion capture markers provide ground truth. 3) 3D position and velocity performance of the policies in Section VI-A.
[…]
@@ V-A Analysis of RMA Module
A property of (1) is that if we allow the drone parameters, $k_{th}$ and $m$, to be variables that can be adjusted online, we can use them to compensate for a wide range of model inaccuracies that are not limited to the thrust and weight of the drone. For simplicity, let $c=\frac{k_{th}}{m}$. Given an additional force vector $\bm{f}_{add}$ in the world frame, to account not only for model inaccuracies within $c$ but also for external forces such as aerodynamic drag and ground effect, we can compute an equivalent $\hat{c}$ in an augmented form of the velocity equation in (1),
$\begin{split}g\bm{z}_{\mathcal{W}}-\hat{c}f_{th}\bm{z}_{\mathcal{B}}&=g\bm{z}_{\mathcal{W}}-cf_{th}\bm{z}_{\mathcal{B}}+\bm{f}_{add}\end{split}.$ (4)
Then, taking the least-squares estimate of $\hat{c}$, we get:
$\min_{\hat{c}}||(\hat{c}-c)f_{th}\bm{z}_{\mathcal{B}}+\bm{f}_{add}||^{2}\;\Rightarrow\;\hat{c}$ $\approx c-\frac{\bm{z}_{\mathcal{B}}^{T}\bm{f}_{add}}{f_{th}}.$ (5)
@@ VI Experiments
In this section, we evaluate our SOUS VIDE policies across three fronts: efficacy of the proposed policy architecture, robustness to dynamic and visual disturbances, and generalization to novel scenarios. We demonstrate that the SV-Net policy, equipped with the RMA module, achieves state-of-the-art performance in zero-shot sim-to-real transfer. We emphasize that in all experiments, the policy does not observe the lateral position $(p_{x},p_{y})$. However, it does observe $p_{z}$ through the onboard time-of-flight sensor input.
We perform all experiments using a quadrotor drone equipped with a PixRacer Pro for low-level body-rate tracking control and an Orin Nano for policy execution, as shown in Fig. 4. The onboard sensing suite consists of an IMU, an ARK Flow sensor, and a monocular camera, with the first two fused via an EKF. The motion capture markers visible in our images and videos are used for diagnostics, enabling trajectory plotting in comparison to the ground truth.
To evaluate performance, we consider four key metrics. Completion: Categorizes trajectories along a discrete spectrum— (✔✔) indicates a fully successful position and orientation tracking with no collisions, (✔) allows for minor collisions with successful recovery, ($\bm{\sim}$) signifies completion of the position component but not the orientation, (✘) denotes failure due to an unrecovered collision, and (✘✘) corresponds to failure due to drifting off-course. Collision Rate (CR): Quantifies the number of collisions per meter traveled. Trajectory Tracking Error (TTE): Measures the $\ell_{2}$-norm of the position error relative to the closest point in the desired trajectory. Finally, Proximity Percentile (PP): Represents the fraction of the trajectory that remains within 30 cm of the intended path. Together, these metrics provide a comprehensive evaluation of trajectory accuracy, robustness, and recovery behavior.
@@ VI-A Policy Architecture Ablations
We evaluated our main policy versus three ablations on a 15-second trajectory that guides the drone through a gate and under a ladder before ending facing a monitor. The desired trajectory is visualized in the GSplat in Fig. 4, along with flights from each policy ablation. All policies were trained on the same expert MPC dataset (180k observation-action pairs) using PyTorch, the Adam optimizer (learning rate 1e-4), for approximately 12 hours on a desktop machine (i9-13900K, RTX 4090, 64GB RAM). The policy ablations are:
@@
  • SV no RMA: A minimal variant comprising only the feature extractor and command network. This serves as our approximation of a zero-shot transfer counterpart to the few-shot transfer described in [17].
@@
Ablation Experiments Completion CR (c/m) TTE (m) PP (%)
✔✔ $\bm{\sim}$✘
SV-Net (ours) 5/5 - - - - 0.00 0.17 96.0
SV no RMA 3/5 - 1/5 - 1/5 0.00 0.65 37.7
SV no pre-train - - 1/5 4/5 - 0.10 0.39 51.4
SV no latent 1/5 2/5 1/5 1/5 - 0.05 0.43 42.0
TABLE I: Table comparing performance of ablations of SV-Net
Figure 5: SV-Net history network’s estimate of $\hat{c}$ with mean $\mu_{\hat{c}}$ overlaid for Section VI-A flights in simulation (left) and real-world (right).
As shown in Table I, SV-Net outperformed all other architectures, achieving a success rate of 100% with no collisions, a TTE of 0.17m and a PP of 96%, more than doubling the performance of SV no RMA.
To study the history network’s performance, we acquired a ground truth estimate of $c=6.03$ by measuring the mass of the drone and recording the throttle command at hover. We found that when pre-trained (SV-Net and SV no latent), the RMA module maintained an estimated $\hat{c}$ value that stayed close to this in both simulation and real-world flights. SV-Net demonstrated the least deviation, with real-world $\mu_{\hat{c}}=6.05,\sigma_{\hat{c}}=0.25$ (illustrated in Fig. 5). In contrast, SV no latent’s estimate is more unstable with $\mu_{\hat{c}}=6.39,\sigma_{\hat{c}}=1.07$ across its five flights. We hypothesize that using the 2nd-to-last layer of the history network improves performance as its higher-dimensional latent code outweighs the minor information loss from skipping the final layer. Consequently, SV no latent suffers from a feedback loop, where poor estimates degrade policy performance, further amplifying estimation errors. We also note that SV no pre-train, which does have a history network but is instead trained directly on control commands, exhibits a highly unstable signal with ($\mu=-2.81,\sigma=9.56$).
Lastly, we observe that using larger datasets, while cheap to synthesize, offers little performance gain while increasing the training time.
@@ VI-B Robustness Experiments
Using the SV-Net result from Section VI-A as a baseline, we conduct five additional experiments, each introducing a distinct disturbance (illustrated in Fig. 6):
@@
  • Dynamic: Four people actively moving within the field of view along the entire trajectory.
  • Static: The gate, ladder, and monitor (present at train time) were removed at runtime while the pillars adjacent to the gate were occluded with white cloth.
  • Payload: A rigid 350g payload ($30\%$ increase in drone weight) was attached below the center-of mass.
@@
Figure 6: Visualization of disturbances and the corresponding position and velocity performance of SV-Net. Lighting: illumination reduced by 60%, Dynamic: human activity in the scene, Static: key objects in training removed at runtime, Payload: 30% increase in mass, Wind: 40 m/s gust from leaf blower. SV-Net maintains adequate performance in all cases.
Robustness Experiments Completion CR (c/m) TTE (m) PP (%)
✔✔ $\bm{\sim}$✘
Baseline 5/5 - - - - 0.00 0.17 96.0
Lighting 4/5 - 1/5 - - 0.00 0.24 64.2
Dynamic 5/5 - - - - 0.00 0.19 86.3
Static - - 4/5 - 1/5 0.06 0.49 25.3
Payload 5/5 - - - - 0.00 0.20 80.1
Wind 5/5 - - - - 0.00 0.20 81.7
TABLE II: Table presenting SV-Net performance under disturbances.
Figure 7: SV-Net estimate of $\hat{c}$ with mean $\mu_{\hat{c}}$ overlaid for Section VI-B flight: payload (left) and wind (right). Wind disturbance region highlighted in green.
The results in Table II show SV-Net consistently demonstrated resilience to dynamic disturbances, payload variations, and wind gusts, maintaining near-baseline performance with minimal impact across all metrics.
When lighting was degraded, the policy struggled to distinguish dark objects from the dim background, particularly near the end, where it drifted away from keeping the (black) monitor in frame. We also tested the policy with less than $40\%$ of the original lumens and the policy consistently failed by drifting off-course from the start location. While the policy handled dynamic scene changes with ease, static changes posed the greatest challenge: i) it underflew waypoints and experienced minor collisions with the occluded pillars, and (ii) it consistently flew through the space where the ladder rungs would have been. Despite these difficulties, the policy reliably tracked the overall trajectory shape, recovered from collisions, and successfully reached the final position in 4 out of 5 flights. These results suggest that the policy is able to retain essential scene information that would otherwise be lost in approaches relying on visual abstractions.
As shown in Fig. 7, the RMA module maintains a stable $\hat{c}$ under wind and payload disturbances, performing nearly identically to the baseline. In the wind disturbance flight, we see a downward spike in $\hat{c}$ that correlates to when the drone passes the leafblower (which is effecting a positive $\bm{f}_{add}$ on the drone). Interestingly, the estimated $\hat{c}$ during the payload flight is perceptibly different from the ground truth estimate updated with the additional mass ($c=4.62$). Given its overall trajectory performance, we believe the RMA module is in fact compensating for inaccuracies in the thrust model in (1), itself a simplified approximation of rotor dynamics.
@@ VI-C Skill-Testing Experiments
In our last set of experiments, we trained three different SV-Net policies, one for each of the following novel scenarios:
  • Multi-Objective: One policy was trained to execute three distinct trajectories within the same scene distinguished by unique objective inputs $O^{k}$. Two trajectories used identical positional splines but traversing in opposite directions, while the third followed a climbing orbit.
@@
  • Cluttered Environment: The policy was deployed close to the ground in a highly cluttered workshop with obstacles spaced as close as 1.0 m apart.
Figure 8: Position and velocity plots for the Multi-Objective (top) and Extended Trajectory (middle) experiments, with the latter’s desired trajectory in its GSplat. We also show a time-lapse of a Cluttered Environment flight (bottom).
Skill-Testing Experiments Completion CR (c/m) TTE (m) PP (%)
✔✔ $\bm{\sim}$✘
Multi-Objective 15/15 - - - - 0.00 0.20 83.4
Extended Trajectory 23/30 4/30 - 3/30 - 0.02 0.24 72.5
Cluttered Environment 13/15 1/15 - - 1/15 0.01 n/a n/a
TABLE III: Table presenting SV-Net performance over novel trajectories.
Figure 9: Quaternion rotation from the IMU during Cluttered Environment flights. The orange outlier trajectory is from the single failed flight.
Visualizations are shown in Fig. 8, with results reported in Table III. The multi-objective policy had mixed success, indicating the need for more robust objective encodings in future work. Though it achieved a $100\%$ collision-free success rate, we observed significant degradation in one of the three tasks, where the drone consistently under-flew the desired trajectory and overshot its end-point In the extended trajectory, the policy performed comparably to the baseline in Section VI-B, with a low CR of 0.02 c/m, TTE of 0.24 m and a PP of 72.5% over 30 attempts. Finally, in the cluttered environment, the policy achieved a $93.3\%$ success rate on a 20 s trajectory through a visually complex scene, demonstrating its robustness in real-world, unstructured environments. Exactly because it is an unstructured environment, there is no motion capture system available for measuring TTE and PP. Instead, we present a time-lapse (bottom of Fig. 8) and the orientation reported by the onboard IMU (Fig. 9).